urlocalgay:

P(A) = 0.23 and P(B) = 0.34. If events A and B are independent, what is the P(A ∩ B)? 0.57 0.11 0.05 0.08

urlocalgay:

I think its D seeing as when I looked it up it said P(A ∩ B) is just P(A) x P(B) and when I multiplied them I got 0.0782 so I assume I'd round that up to 0.8

toga:

If events A and B are independent, then the probability of A and B happening together (i.e., the intersection of A and B) is equal to the product of their individual probabilities, which is: P(A ∩ B) = P(A) × P(B) Substituting the given probabilities, we have: P(A ∩ B) = 0.23 × 0.34 = 0.0782 Rounding to two decimal places, we get: P(A ∩ B) ≈ 0.08 Therefore, the answer is (d) 0.08.

urlocalgay:

Thank you!

toga:

@urlocalgay wrote:
Thank you!
np